OpenStudy (anonymous):

the P.E of body at height of 3m from the ground is 30J.its K.E AT HEIGHT OF 1m will be (g=9.8m/s^2)

OpenStudy (cleonideu):

Is this a question?

OpenStudy (anonymous):

yes :)

OpenStudy (cleonideu):

Hmmm

OpenStudy (cleonideu):

*puts on thinking cap*

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

\[E_i = E_f\] \[PE_I + KE_i = PE_f + KE_f\] \[PE_I +\cancel{ KE_i} = PE_f + KE_f\] \[PE = mgh\] \[ PE_{3m} = mg(3m)\] \[PE_{3m}=PE_{1m}+KE_{1m}\] \[KE_{1m} = PE_{3m}-PE_{1m} \]

OpenStudy (anonymous):

30 - 10 = 20 J .am i right @AllTehMaffs

OpenStudy (anonymous):

and y we cancelled initial K.E ?

OpenStudy (anonymous):

Kinetic energy only applies if the object is moving; the problem doesn't say that is is, so I think we can assume that it starts at rest. And you're right :) How did you get that answer though?

OpenStudy (anonymous):

P.E = mgh

OpenStudy (anonymous):